the length of a field is 3 times its width if the perimeter is 64m what is the width of the field?
Answer by waynest(281)

length = 3w
p = 2L + 2w
64 = 2(3w) + 2w
64 = 6w + 2w
64 = 8w
64/8 = 8w/8
8 = w
check:
64 = 2(3(8)) + 2(8)
64 = 2(24) + 16
64 + 48 + 16
64 = 64